Explain the key role of a central bank (such as the Federal Reserve) in the monetary system.

What happens to the money supply when a central bank (such as the Federal Reserve) buys bonds? Explain.

You run a bank. The current reserve ratio mandates holding reserves equal to 20% of deposits. If someone comes into your bank and deposits $10,000, by how much will the money supply in the economy increase?

You have equity (a capital share) in a bank. The bank has reserves, loans, and securities as well as deposits and capital. The bank suffers a sudden loss of value in one of its securities (assets). What will happen to the amount of capital the bank has and the leverage ratio?
Answer #1

The key role is to stabilize economy, reduce unemployment and prevent inflation or depression..in other words it counters bad effects of trade cycle and thus promotes growth

2 it buys bonds in lieu of money. Thus public recieves more cash.Due to credit creation process money supply increases by quite large amount

3 money multiplier =1/RR=1/20%=5.so increase =5(10,000)=50,000

4 Amount of capital will remain same but assets will fall. Leverage ratio will thus ruse